OpenStudy (anonymous):

2x+y=10 -5x-y=-19 Solve using system of equations

OpenStudy (kidrah69):

this is a system of equations lets use substitution :) \[-5x-(-2x+10)=-19 \] i subtracted 2x from the first equation to get y equals.... and plugged it into the 2nd equation can you solve from there?

OpenStudy (anonymous):

Wait woah aren't we supposed to solve for x first?

OpenStudy (kidrah69):

this is the way to do it :D solve the equation

OpenStudy (anonymous):

I got x=3

OpenStudy (anonymous):

@kidrah69

OpenStudy (anonymous):

And i got y=4

OpenStudy (imtiaz7):

lets do: multiply both equations by 2 we get: 4x+2y=20 and -10x-2y=-38 now when we add them we got: -6x=-18 x=3 now put it in first equation we get: y=4

OpenStudy (imtiaz7):

Yup you are right!

OpenStudy (anonymous):

I got another one

OpenStudy (imtiaz7):

ok am here :)

OpenStudy (anonymous):

5x+4y=-2 3x+4y=10

OpenStudy (imtiaz7):

now post this queation again plz